### v2

# The Baryonic Tully-Fisher Relation of HI-bearing Low Surface Brightness Galaxies Implies Their Formation Mechanism

Abstract
We investigate the baryonic Tully-Fisher relation in low surface brightness galaxies selected from the Arecibo Legacy Fast ALFA survey. We find that the $\rm HI$-bearing low surface brightness galaxies still follow the baryonic Tully-Fisher relation of typical late-type galaxies, with a slope of approximately 4 in the baryonic mass versus rotational velocity diagram on the logarithmic scale, i.e., $M_{\rm{b}}\propto v_{\rm{rot}}^{4}$. Our findings suggest that the matter distributions in low surface brightness galaxies may resemble that of general late-type galaxies, and hint that low surface brightness galaxies may not originate from dark matter halos of low densities or stronger/weaker feedback processes, but may emerge from dark matter halos with high spin values.
keywords:
galaxies: kinematics and dynamics – galaxies: formation – galaxies: evolution
pubyear: 2024pagerange: The Baryonic Tully-Fisher Relation of HI-bearing Low Surface Brightness Galaxies Implies Their Formation MechanismThe Baryonic Tully-Fisher Relation of HI-bearing Low Surface Brightness Galaxies Implies Their Formation Mechanism
1 Introduction
Low surface brightness galaxies (LSBGs) constitute a special population of galaxies with central surface brightness at least one magnitude fainter than that of the sky background (i.e., $B$-band surface brightness $\rm\geq 22.5\ mag\cdot arcsec^{-2}$; Impey & Bothun, 1997; Bothun et al., 1997). In the fields, LSBGs have been observed to exhibit higher HI fractions (Du et al., 2015; He et al., 2020), lower star formation rates (Wyder et al., 2009; Rong et al., 2020b), and low metallicities (Kuzio de Naray et al., 2004). Also, LSBGs tend to appear in low-density environments (Pérez-Montaño & Cervantes Sodi, 2019; Mo et al., 1994; Galaz et al., 2011).
The formation and evolution mechanisms of LSBGs, which play a crucial role in refining our current models of galaxy formation and evolution, have gathered much attention ever since their initial detection. Early investigations often regarded LSBGs as galaxies formed in low-density dark matter halos (Dekel & Silk, 1986; McGaugh, 1992; Mo et al., 1994). However, more recent studies have presented more alternative scenarios. One of the most classical theories of the formation mechanism of LSBGs is that they are galaxies hosted in dark matter halos belonging to the high-spin tail of the spin distribution. In theoretical works and simulations, the halos of LSBGs are found to have higher spin parameters compared with high surface brightness galaxies, regardless of whether they are relatively massive (Mo et al., 1998; Kim & Lee, 2013; Kulier et al., 2020; Pérez-Montaño et al., 2022), or belong to ultra-diffuse galaxies (UDGs, Rong et al., 2017; Amorisco & Loeb, 2016; Rong et al., 2024) – a specific subset of LSBGs of smaller masses. Additionally, some studies suggest that low-mass LSBGs may represent failed $L^{\rm*}$ galaxies (van Dokkum et al., 2015a, b), or originate from early co-planar mergers (Wright et al., 2021) or stellar feedback processes (Chan et al., 2018; Di Cintio et al., 2019), or birth during tidal interactions (Rong et al., 2020a; Carleton et al., 2019), or form through multiple pathways (Papastergis et al., 2017). High-mass LSBGs, alternatively, may form through mergers (Saburova et al., 2018; Di Cintio et al., 2019; Zhu et al., 2023), two-stage process with external gas accretions (Saburova et al., 2021), or dynamical evolution driven by bars (Noguchi, 2001).
Nevertheless, despite extensive efforts, our present comprehension of the formation and evolution of LSBGs is still limited. One of the key reasons is that studies on the mass distributions and baryonic fractions in LSBGs have not achieved a consensus, and a comprehensive scenario has not been established to constrain the current formation and evolution models of LSBGs. Many studies have shown that the dynamics of LSBGs may be dominated by dark matter (e.g. de Blok & McGaugh, 1997; Mowla et al., 2017), yet UDGs may exhibit a deficiency in dark matter content (e.g. van Dokkum et al., 2018; Mancera Piña et al., 2019). As for the stellar-to-halo mass relation ($M_{\rm*}-M_{\rm halo}$), while Prole et al. (2019) propose that LSBGs form a continuous extension of typical dwarfs in the $M_{\rm*}-M_{\rm halo}$ diagram, other studies reveal significant deviations for UDGs from the typical $M_{\rm*}-M_{\rm halo}$ relation (e.g. van Dokkum et al., 2018). Additionally, some giant LSBGs appear to be dominated by baryons at their central regions, akin to the high surface brightness galaxy (HSBG) counterparts (Lelli et al., 2010); the stellar-to-halo mass relation of the giant LSBGs and HSBGs also exhibits a similarity. All of these confusing signs of diversity in matter distribution among LSBGs could offer valuable insights into their varied formation and evolution pathways. Consequently, a comprehensive understanding of the matter distributions in LSBGs is essential for refining our current models of galaxy formation and evolution.
The baryonic Tully-Fisher relation (BTFR) is a correlation between the rotational velocity, $v_{\rm rot}$, and the baryonic mass, $M_{\rm b}$, of galaxies. Over the past two decades, many studies have shown that late-type galaxies and dwarf galaxies consistently adhere to a power-law form of BTFR, despite their wide ranges of baryonic mass, luminosity, and size. BTFR could be written as,
$M_{\rm b}\propto v_{\rm rot}^{\beta},$ (1)
where the logarithmic slope, $\beta$, typically falls within a range from 3 to 4 (e.g. McGaugh et al., 2000; Begum et al., 2008; Papastergis et al., 2016; Lelli et al., 2016b; Sales et al., 2017; Goddy et al., 2021). Notably, some studies have also discovered that early-type galaxies may also follow the BTFR of late-type galaxies (den Heijer et al., 2015; Serra et al., 2012).
In the context of $\rm\Lambda CDM$ paradigm, the presence of BTFR is considered as a correlation between the virial velocity of the dark matter halo and the baryonic mass enclosed within that halo (McGaugh, 2012; Mo et al., 1998). Thus, it can serve as a valuable tool to probe the distribution of baryonic matter and dark matter within galaxies. Furthermore, simulations have highlighted the significance of feedback processes in shaping the observed BTFR (e.g. Governato et al., 2010; Piontek & Steinmetz, 2011; Dutton, 2012), suggesting that investigating the BTFR can serve as a valuable means to constrain the feedback mechanisms at work within galaxies. Additionally, BTFR is highly related to other dynamical relationships such as the radial acceleration relation (e.g. McGaugh et al., 2016; Lelli et al., 2017). Exploring the BTFR could also provide insights into new gravitational theories, as a BTFR with $\beta=4$ is well predicted by the MOdified Newtonian Dynamics (MOND, Milgrom, 1983). Given its multifaceted implications, the BTFR remains one of the most extensively studied dynamical relations in astrophysics.
As anticipated, recent studies examining the BTFR of LSBGs also show some potential signs of tension. The previous studies of, for instance, Lelli et al. (2016b) (see also Zwaan et al., 1995), have explored that the residuals of BTFR do not change with galactic surface brightness, while some of the recent investigations (Hu et al., 2023; Karunakaran et al., 2020; Rong et al., 2024) reveal that UDGs may significantly deviate from the BTFR of typical late-type galaxies. Therefore, a thorough and meticulous examination of the BTFR for LSBGs becomes essential. Such an analysis may not only indicate whether the matter distribution within LSBGs aligns with that of typical late-type and dwarf galaxies, but also provide valuable constraints on the formation mechanisms of LSBGs.
In this work, we will first introduce our sample selection in Section 2, and then describe our methods for calculating $M_{\rm b}$ and $v_{\rm rot}$ in Section 3. We will then present our results in Section 4, and finally discuss the potential selection biases, the radial acceleration relation, and the possible formation mechanism of LSBGs in Section 5. The summary is in Section 6.
2 Sample Selection
[…]
From the parent sample, we first remove LSBGs with $g$-band apparent axis ratios $b/a>0.72$, approximately corresponding to the threshold of inclination angles of $i\lesssim 45^{\circ}$, because that the face-on galaxies may introduce large uncertainties in the measurements of the rotation velocities (see equation (10)). We then exclude LSBGs with low $\rm HI$ spectral signal-to-noise ratios ${\rm SNR}<10$, and visually remove LSBGs with suspicious HI spectra.
Due to the bad spatial resolution of ALFALFA (Durbala et al., 2020, being $3.8^{\prime}\times 3.3^{\prime}$;), if a galaxy has a close companion galaxy, its $\rm HI$ spectrum may be contaminated by the companion. To eliminate this possibility, we remove the sample galaxies with neighbouring galaxies within the projected radii of $3.8^{\prime}$ (i.e., approximately the beam size of Arecibo) and radial velocity differences of $\Delta v<500{\rm km/s}$. The redshift of a LSBG is estimated from the central velocity of its $\rm HI$ spectrum, while the redshift of the neighbouring galaxies comes from the SpecObj and Photoz databases of SDSS. Finally, 204 LSBGs are left in our sample.
We also calculate the kurtosis coefficients, $k_{\rm 4}$, of the HI spectra (Papastergis et al., 2016; El-Badry et al., 2018) for our sample galaxies, as
$k_{4}=M_{4}/(\sigma^{2})^{2}-3,$ (2)
where $M_{4}$ and $\sigma^{2}$ are calculated as,
[…]
In equations (3) and (4), $\nu$ and $F(\nu)$ represent the frequency and the flux density, respectively, and $\bar{\nu}$ is defined as,
[…]
Here, $\nu_{\rm min}$ and $\nu_{\rm max}$ correspond to the minimal and maximum frequencies between which $W_{20}$ (see Section 3 for the definition of $W_{20}$) is measured, respectively. The coefficient $k_{\rm 4}$ quantifies the “peakiness” of a spectrum and thus can indicate whether a spectrum is single-peaked or double-horned. The single-peaked galaxies may have kinematics strongly influenced by velocity dispersion (El-Badry et al., 2018), or HI disks that are not extended enough to trace the flat part of the rotation curve (Papastergis et al., 2016). As a consequence, their HI-widths might not represent their rotation velocities (McGaugh, 2012; Verheijen, 1997). Therefore, we perform a cut at the HI spectrum, $k_{\rm 4}<-1$, in order to remove the single-peaked galaxies.
[…]
3 Method
We study the BTFRs of LSBGs and HSBGs by directly estimating the baryonic masses, $M_{\rm b}$, and rotation velocities, $v_{\rm rot}$, of the galaxies in the two samples.
The baryonic mass of a galaxy in our samples is mainly composed of the mass of stars and the mass of cold gas. The stellar mass $M_{\rm*}$ can be estimated with the galaxy luminosity in the $W1$-band (denoted as $L_{W1}$) from the Wide-field Infrared Survey Explorer (Wright et al., 2010), using the method proposed by McGaugh & Schombert (2015), as
$M_{\rm*}/M_{\rm\odot}=0.45L_{W1}/L_{\rm\odot}.$ (6)
As for the mass of the cold gas, $M_{\rm g}$, we estimate it by adding up the masses of HI and helium,
$M_{\rm g}=1.33M_{\rm HI},$ (7)
where $M_{\rm HI}$ is the HI mass. Equation (7) will certainly underestimate the mass of the gas since it neglects the mass of the molecular gas and hot ionized gas. However, the molecular gas fraction in the total gas component is low and thus negligible (McGaugh, 2012). Also, the hot ionized gas usually resides outside the HI disk where our rotational velocity is measured (Dutton, 2012), so it can be neglected. Therefore, approximately we estimate the baryonic mass of a galaxy as,
$M_{\rm b}=M_{\rm*}+M_{\rm g}.$ (8)
For the rotation velocity, we estimate it using the $20\%$ peak width of a HI spectrum (Guo et al., 2020; Hu et al., 2023; Rong et al., 2024, $W_{\rm 20}$;), as,
$v_{\rm rot}=\frac{W_{20}}{2\sin i},$ (9)
where $i$ is the inclination angle. $W_{20}$ employed in equation (9) has been corrected for instrumental broadening and redshift (Guo et al., 2020; Kent et al., 2008). In equation (9), we assume an alignment between the inclination angle of the stellar disk and HI disk in a galaxy, and then use the inclination angle of the stellar disk to estimate the rotation velocity. The inclination is estimated from the apparent axis ratio, as (e.g. Guo et al., 2020; Hu et al., 2023; Du et al., 2019; Rong et al., 2024; Begum et al., 2008),
$\sin i$ $=\sqrt{\frac{1-(b/a)^{2}}{1-q_{0}^{2}}},$ (10)
where $q_{0}$ denotes the intrinsic axis ratio seen edge-on. We take $q_{0}=0.2$, which is commonly used in the previous studies (e.g. Du et al., 2019; Guo et al., 2020; Giovanelli et al., 1997; Tully et al., 2009). For those edge-on galaxies with $g$-band $b/a\leq 0.2$, we set $\sin i=1$. Considering that the stellar disk and gas disk in galaxies may not be perfectly co-planar and often exhibit a small inclination difference of ${\rm\delta}i<20^{\circ}$ (e.g. Gault et al., 2021), we assume that the misalignment angle between the gas disk and the stellar disk, ${\rm\delta}i$, follows a Gaussian distribution centred at $0^{\circ}$ with a standard deviation of $\sigma=20^{\circ}$. We treat $\sigma=20^{\circ}$ as the uncertainty associated with $i$ in our study.
4 Results
In Figure 1, we observe that the distributions of the LSBGs and HSBGs exhibit no significant deviation in the $v_{\rm rot}-M_{\rm b}$ diagram. Subsequently, we impose a power-law equation,
$\log(M_{\rm b})=\alpha+\beta\log(v_{\rm rot}),$ (11)
where $\alpha$ and $\beta$ are both free parameters, to fit the BTFRs of the HSBGs and LSBGs, respectively. The best-fitting parameters of the BTFR of HSBGs are $\alpha_{\rm HSBG}=1.83\pm 0.51$ and $\beta_{\rm HSBG}=3.91\pm 0.24$, while the best-fitting parameters of LSBGs are $\alpha_{\rm LSBG}=-1.45\pm 1.49$ and $\beta_{\rm LSBG}=5.67\pm 0.78$.
We further split LSBGs and HSBGs into the different $\log v_{\rm{rot}}$ bins. Within each bin, we calculate the median value of $\log M_{\rm{b}}$ and the corresponding $1\sigma$ standard deviation for both samples. As shown in Figure 1, the binned points of LSBGs (depicted by the blue squares), as well as the binned points of HSBGs (depicted as the black squares), comfortably lie within the $1\sigma$ uncertainty region of the tightest BTFR of typical late-type galaxies reported by Lelli et al. (2019) (the red-solid line and the red-shaded region denote the best-fitting BTFR of typical late-type galaxies and its $1\sigma$ uncertainty, respectively, as shown in Fig. 1), with $\beta\simeq 3.85\pm 0.09$.
Therefore, we conclude that the BTFRs of LSBGs and HSBGs show no significant deviation. Both galaxy samples follow the BTFR of typical late-type galaxies, with a BTFR slope of approximately 4.
Figure 1: $v_{\rm rot}-M_{\rm b}$ diagram for our samples. The green circles represent LSBGs, while the grey circles represent HSBGs, and the orange circles are SPARC galaxies from Lelli et al. (2019). All the error bars denote $\rm 1\sigma$ uncertainties. The blue squares represent the median values of LSBGs while the black squares represent that of HSBGs. The red-solid line is the tightest BTFR reported by Lelli et al. (2019). The red-shaded area is the $\rm 1\sigma$ intrinsic scatter of the red-solid line, which is considered as the $\rm 1\sigma$ uncertainty of the tightest BTFR. The sub-panels display the number-density distribution of LSBGs (green histograms) and HSBGs (grey histograms).
5 Discussion
5.1 Selection biases
In our study, the SNR threshold used to select the LSBG and HSBG samples may exclude some galaxies with similar HI masses but wider HI line widths from our parent sample, which might potentially introduce deviations in the slope of BTFR. To investigate the possible selection effects, we adjust the SNR threshold by increasing it from 10 to 15, or 20.
We find that, for HSBGs, the best-fitting $\beta$ of their BTFR changes to $4.05\pm 0.21$ for SNR $>15$, and $4.05\pm 0.26$ for SNR $>20$, still in good agreement with the best-fitting $\beta$ value for SNR$>10$. For LSBGs, the best-fitting $\beta$ becomes $\beta\sim 6.04\pm 0.66$ (SNR $>15$) and $\sim 6.13\pm 0.82$ (SNR $>20$), which is also similar to the $\beta\sim 5.67\pm 0.78$ when applying the selection criterion of SNR$>10$, considering the large uncertainties of the $\beta$ values.
Additionally, the binned points of LSBGs and HSBGs within the different $\log v_{\rm{rot}}$ for the different SNR thresholds are always located within the $1\sigma$ uncertainty region of the BTFR of typical late-type galaxies. Consequently, we conclude that, for both of LSBG and HSBG samples, the selection effect introduced by the SNR threshold is negligible.
5.2 The radial acceleration relation
Apart from BTFR, recently, a series of studies (McGaugh et al., 2016; Lelli et al., 2017; Li et al., 2018; Richards et al., 2018; Rong et al., 2018; Chae et al., 2020) have revealed another tight correlation known as the radial acceleration relation (RAR). RAR indicates that the acceleration $g_{\rm bar}$ produced by the baryonic mass and the acceleration $g_{\rm obs}$ produced by the dynamical mass at any radius in many types of galaxies consistently follow a correlation of
$g_{\rm obs}=\frac{g_{\rm bar}}{1-e^{-\sqrt{g_{\rm bar}/a_{0}}}},$ (12)
where $a_{0}\simeq(1.20\pm 0.02\pm 0.24)\times 10^{-10}\ {\rm m/s^{2}}$ (McGaugh et al., 2016; Lelli et al., 2017) is constant. In the $\rm\Lambda CDM$ paradigm, RAR plausibly suggests a universally established link between the mass distributions of baryonic matter and dark matter in galaxies.
We estimate $g_{\rm obs}$ and $g_{\rm bar}$ of each galaxy at its HI-radius ($R_{\rm HI}$) using the following equations,
$g_{\rm obs}=\frac{v_{\rm rot}^{2}}{R_{\rm HI}},$ (13)
$g_{\rm bar}=x_{\rm*}\frac{GM_{\rm*}}{R^{2}_{\rm HI}}+x_{\rm g}\frac{GM_{\rm g}}{R^{2}_{\rm HI}},$ (14)
where $G$ is the gravitational constant. $x_{\rm*}$ and $x_{\rm g}$ in equation (14) are correction factors assuming that the cold gas and stars are distributed in two razor-thin disks, respectively. They are calculated as (Freeman, 1970),
$x_{\rm*}={\rm I_{0}}(\frac{R_{\rm HI}}{2R_{d,*}}){\rm K_{0}}(\frac{R_{\rm HI}}{2R_{d,*}})-{\rm I_{1}}(\frac{R_{\rm HI}}{2R_{d,*}}){\rm K_{1}}(\frac{R_{\rm HI}}{2R_{d,*}}),$ (15)
$x_{\rm g}={\rm I_{0}}(\frac{R_{\rm HI}}{2R_{d,g}}){\rm K_{0}}(\frac{R_{\rm HI}}{2R_{d,g}})-{\rm I_{1}}(\frac{R_{\rm HI}}{2R_{d,g}}){\rm K_{1}}(\frac{R_{\rm HI}}{2R_{d,g}}),$ (16)
where $\rm I_{n}$ and $\rm K_{n}$ are the modified Bessel functions of the first and second kind, respectively; $R_{\rm d,*}$ and $R_{\rm g,*}$ involved are the scale lengths of the stellar and HI disks which can be derived by using the same method described in the work of Rong et al. (2024).
$R_{\rm HI}$ corresponds to the outer region of a galaxy where the surface density of HI equals $1\ M_{\rm\odot}/{\rm pc^{2}}$ (Broeils & Rhee, 1997). $R_{\rm HI}$ can be estimated from the following tight correlation between HI mass and $R_{\rm HI}$, which has been observed in various types of galaxies including LSBGs (Wang et al., 2016; Gault et al., 2021; Lutz et al., 2018),
$\log(\frac{2R_{\rm HI}}{\rm kpc})=(0.506\pm 0.003)\log(\frac{M_{\rm HI}}{M_{\rm\odot}})-(3.293\pm 0.009).$ (17)
In Figure 2, at $R_{\rm HI}$, both our sample LSBGs and HSBGs follow the universal RAR (McGaugh et al., 2016; Lelli et al., 2017, as illustrated by the red components in Fig. 2). The similarity of RARs between LSBG and HSBGs may suggest comparable mass distributions in the two galaxy samples.
It is worth noting that, at a large radius of a galaxy where $g_{\rm bar}\ll a_{0}$, such as $R_{\rm{HI}}$, the RAR described by equation (12) is actually equivalent to the BTFR with $\beta=4$ (Lelli et al., 2017).
Figure 2: The RAR of our sample galaxies, overplotted by data from Lelli et al. (2016a). The circles and squares are the same as those in Figure 1. The red-solid line is the best-fitted RAR in previous studies (McGaugh et al., 2016; Lelli et al., 2017), and the grey-dot-dashed line is the identical line.
5.3 Implications on the formation mechanisms of LSBGs
Previous theoretical studies posit that LSBGs may form in low-density dark matter halos (e.g. Dekel & Silk, 1986; McGaugh, 1992). However, this model predicts a significant difference in the BTFRs between LSBGs and HSBGs (McGaugh, 2021), which contradicts our findings. Therefore, the formation model based on low-density dark matter halos may not refer to LSBGs.
Many observational studies also suggest that LSBGs might be dominated by dark matter compared with HSBGs (e.g. de Blok & McGaugh, 1997). Therefore, we delve deeper into understanding the fraction of dark matter in LSBGs. Note that BTFR could be expressed as  [^1] Equation (19) is from the lecture on Tully-Fisher relation; see also Aaronson et al. (1979) and Zwaan et al. (1995).,
$M_{\rm b}\simeq\frac{f_{\rm b}^{2}(R)}{Gg_{\rm bar}(R)}v_{\rm rot}^{4}.$ (18)
where $R$ corresponds to a large radius in the outer region of a galaxy (such as $R_{\rm{HI}}$), and
$f_{\rm b}(R)=\frac{M_{\rm b}}{M_{\rm b}+M_{\rm DM}(R)},$ (19)
is the baryonic mass fraction within $R$, where $M_{\rm{DM}}(R)$ denotes the dark matter mass enclosed within $R$. $g_{\rm bar}(R)$ is the baryonic acceleration at $R$, which could be calculated with equation (14). Consequently, $f_{\rm b}(R_{\rm{HI}})$ of the LSBGs and HSBGs can be estimated with equation (18).
We examine the variation of $f_{\rm b}(R_{\rm HI})$ as a function of $v_{\rm{rot}}$ for both LSBGs and HSBGs. As shown in Fig. 3, in statistics, $f_{\rm b}(R_{\rm HI})$ increases with rising $v_{\rm{rot}}$, with a slope of $\sim 1.02$ using a bisector fitting (Isobe et al., 1990). Since the dark matter halo mass of a galaxy is tightly related to $v_{\rm{rot}}$, as $M_{\rm{h}}\propto v_{\rm{rot}}^{3}$ (Mo et al., 1998), the correlation between $f_{\rm b}(R_{\rm HI})$ and $v_{\rm{rot}}$ hints at a positive association between the baryonic mass fractions and dark matter halo masses for these galaxies (Ayromlou et al., 2023; Duffy et al., 2010). This correlation does not exceed our expectations, given that our galaxy samples are dominated by the low-mass galaxies. Halos with lower $M_{\rm{h}}$ have shallower potentials (Duffy et al., 2010), making it more difficult to prevent the galactic interstellar medium (ISM) against expulsion by the feedback processes in these shallower potentials, which leads to lower $f_{\rm b}(R_{\rm{HI}})$ (e.g. Bryan et al., 2013; Duffy et al., 2010).
Subsequently, in order to compare the $f_{\rm b}(R_{\rm{HI}})$ of the LSBG and HSBG samples with similar dark matter halo masses, we further focus on a specific $v_{\rm{rot}}$ bin defined by $\log v_{\rm rot}\in[1.87,2.05]$ (as depicted by the blue-shaded region in Fig. 3). We have 52 LSBGs and 52 HSBGs in this small bin, exhibiting similar distributions of $v_{\rm rot}$ with median $\log v_{\rm rot}$ values of $1.94\pm 0.05$ (for LSBGs) and $1.97\pm 0.05$ (for HSBGs). The Kuiper test between the distributions of $v_{\rm rot}$ of LSBGs and HSBGs in this small bin also gives a large $p$-value of $0.77$, indicating no significant difference in $v_{\rm rot}$. We then compare the $f_{\rm b}(R_{\rm HI})$ distributions of the LSBG and HSBG sample galaxies within this $v_{\rm rot}$ bin. The median values of the $f_{\rm b}(R_{\rm HI})$ distributions of LSBGs and HSBGs are $-0.94\pm 0.15$ and $-0.91\pm 0.30$, respectively; the Kuiper test for the two $f_{\rm b}(R_{\rm HI})$ distributions also yields a large $p$-value of $0.34$.
Therefore, we note that, for the LSBGs and HSBGs with similar halo masses, their baryonic fractions are also comparable. LSBGs are not particularly dark matter dominated galaxies, compared with the HSBG counterparts. Furthermore, this result also plausibly suggests comparable feedback power among the LSBGs and HSBGs with similar halo masses (i.e., similar gravitational potentials). Consequently, we conclude that feedback may not be the primary driver of the different surface brightness between these HI-bearing galaxies. In other words, these HI-bearing LSBGs may not originate from stronger or weaker feedback.
Alternatively, our finding aligns with the simulation results where the LSBGs can be well reproduced by higher dark matter halo spins (Pérez-Montaño et al., 2022). Therefore, we conclude that our result may lend support to the theorem that LSBGs may originate from high-spin dark matter halos (Mo et al., 1998).
Figure 3: The $v_{\rm rot}-f_{\rm b}(R_{\rm HI})$ relation of our sample galaxies. The circles are the same as those in Figure 1. The blue-shaded area represents the bin in which we perform Kuiper tests in Section 5.3. The black-dashed line is the best-fitted line for the data points.
5.4 Implications on MOND
Except for the aforementioned formation models of LSBGs in the framework of $\rm\Lambda CDM$, MOND may also offer a reasonable explanation for the BTFR and RAR of LSBGs (e.g. Milgrom, 1983; Wittenburg et al., 2020).
As demonstrated in Section 4 and 5.2, both LSBGs and HSBGs adhere to a BTFR with a slope indistinguishable from 4, and may follow the same RAR as given by equation (12). These results adhere to the theoretical predictions by MOND (Milgrom, 1983). Therefore, our result cannot exclude the MOND theory.
6 Summary
Taking advantage of the ALFALFA HI survey data, we have selected the HI-bearing LSBGs and HSBGs, and estimated their baryonic masses $M_{\rm b}$ and rotation velocities $v_{\rm rot}$. Our findings reveal that, statistically, the HI-bearing LSBGs and HSBGs also conform to the BTFR found in typical late-type galaxies as well as ordinary dwarf galaxies with slope of $\beta\sim 4$. These results hint that LSBGs may owe their properties to the high spins of their host dark matter halos, rather than the halo densities or feedback processes. Additionally, our findings do not rule out the Modified Newtonian Dynamics (MOND) theory.
Our findings conflict with the BTFR results of UDGs (Karunakaran et al., 2020; Hu et al., 2023; Rong et al., 2024). This is primarily because that our LSBG sample contains very few UDGs. Indeed, almost all of our LSBGs are classic LSBGs, rather than UDGs with extremely large effective radii. Therefore, we propose that UDGs may be a distinct galaxy population with different BTFR and mass distributions, compared with typical LSBGs. They may also have distinct formation mechanisms that warrant further investigations.
Acknowledgements
We sincerely thank Xufen Wu, Haochen Jiang, Jingshuo Yang, Shijiang Chen and Yu Chen for the inspiring discussions during this study. Y.R. acknowledges supports from the NSFC grant 12273037, and the CAS Pioneer Hundred Talents Program (Category B), as well as the USTC Research Funds of the Double First-Class Initiative (No. YD2030002013). H.H. is supported by the Fundamental Research Funds for the Central Universities, the CAS Project for Young Scientists in Basic Research Grant No. YSBR-062, the National SKA Program of China No. 2022SKA0110201, and the NSFC grant No. 12033008.
Data Availability
Data are available if requested.

### v3

@@
 Low surface brightness galaxies (LSBGs) constitute a special population of galaxies with central surface brightness at least one magnitude fainter than that of the sky background (i.e., $B$-band surface brightness $\rm\geq 22.5\ mag\cdot arcsec^{-2}$; Impey & Bothun, 1997; Bothun et al., 1997).
-In the fields, LSBGs have been observed to exhibit higher HI fractions (Du et al., 2015; He et al., 2020), lower star formation rates (Wyder et al., 2009; Rong et al., 2020b), and low metallicities (Kuzio de Naray et al., 2004).
-Also, LSBGs tend to appear in low-density environments (Pérez-Montaño & Cervantes Sodi, 2019; Mo et al., 1994; Galaz et al., 2011).
-The formation and evolution mechanisms of LSBGs, which play a crucial role in refining our current models of galaxy formation and evolution, have gathered much attention ever since their initial detection.
-Early investigations often regarded LSBGs as galaxies formed in low-density dark matter halos (Dekel & Silk, 1986; McGaugh, 1992; Mo et al., 1994).
-However, more recent studies have presented more alternative scenarios.
-One of the most classical theories of the formation mechanism of LSBGs is that they are galaxies hosted in dark matter halos belonging to the high-spin tail of the spin distribution.
-In theoretical works and simulations, the halos of LSBGs are found to have higher spin parameters compared with high surface brightness galaxies, regardless of whether they are relatively massive (Mo et al., 1998; Kim & Lee, 2013; Kulier et al., 2020; Pérez-Montaño et al., 2022), or belong to ultra-diffuse galaxies (UDGs, Rong et al., 2017; Amorisco & Loeb, 2016; Rong et al., 2024) – a specific subset of LSBGs of smaller masses.
+Many isolated LSBGs have been observed to exhibit higher HI fractions (e.g., Catinella et al., 2018; Schombert & McGaugh, 2021; Du et al., 2015; He et al., 2020).
+However, some studies (e.g., Martin et al., 2019; Galaz et al., 2002) alternatively found that some LSBGs exhibit the same or even smaller gas fractions compared to high surface brightness galaxies (HSBGs).
+LSBGs usually have low star formation rates (Wyder et al., 2009; Rong et al., 2020b), low metallicities (Kuzio de Naray et al., 2004; Schombert & McGaugh, 2021), and tend to appear in low-density environment (Pérez-Montaño & Cervantes Sodi, 2019; Mo et al., 1994; Galaz et al., 2011).
+The formation mechanism of LSBGs, which plays a crucial role in refining our empirical galaxy formation models, have gathered much attention ever since their initial detection.
+Early studies often regarded LSBGs as galaxies formed in low-density dark matter halos (Dekel & Silk, 1986; McGaugh, 1992; Mo et al., 1994).
+However, recent studies have presented more alternative scenarios.
+One classical formation scenario is that LSBGs are galaxies hosted in dark matter halos with higher halo spin.
+In cosmological simulations, the halos of LSBGs are found to have higher spin parameters compared with HSBGs, regardless of whether they are relatively massive (Mo et al., 1998; Kim & Lee, 2013; Kulier et al., 2020; Pérez-Montaño et al., 2022), or belong to ultra-diffuse galaxies (UDGs, Rong et al., 2017; Amorisco & Loeb, 2016; Rong et al., 2024) – a specific subset of LSBGs of smaller masses (e.g., van Dokkum et al., 2015a, 2018; Mancera Piña et al., 2018, 2019).
 Additionally, some studies suggest that low-mass LSBGs may represent failed $L^{\rm*}$ galaxies (van Dokkum et al., 2015a, b), or originate from early co-planar mergers (Wright et al., 2021) or stellar feedback processes (Chan et al., 2018; Di Cintio et al., 2019), or birth during tidal interactions (Rong et al., 2020a; Carleton et al., 2019), or form through multiple pathways (Papastergis et al., 2017).
@@
 One of the key reasons is that studies on the mass distributions and baryonic fractions in LSBGs have not achieved a consensus, and a comprehensive scenario has not been established to constrain the current formation and evolution models of LSBGs.
-Many studies have shown that the dynamics of LSBGs may be dominated by dark matter (e.g.
-de Blok & McGaugh, 1997; Mowla et al., 2017), yet UDGs may exhibit a deficiency in dark matter content (e.g.
-van Dokkum et al., 2018; Mancera Piña et al., 2019).
-As for the stellar-to-halo mass relation ($M_{\rm*}-M_{\rm halo}$), while Prole et al.
-(2019) propose that LSBGs form a continuous extension of typical dwarfs in the $M_{\rm*}-M_{\rm halo}$ diagram, other studies reveal significant deviations for UDGs from the typical $M_{\rm*}-M_{\rm halo}$ relation (e.g.
-van Dokkum et al., 2018).
-Additionally, some giant LSBGs appear to be dominated by baryons at their central regions, akin to the high surface brightness galaxy (HSBG) counterparts (Lelli et al., 2010); the stellar-to-halo mass relation of the giant LSBGs and HSBGs also exhibits a similarity.
+Many studies have shown that the dynamics of LSBGs may be dominated by dark matter (e.g., de Blok & McGaugh, 1997; Mowla et al., 2017; Swaters et al., 2003; Read et al., 2017; Mancera Piña et al., 2022), yet UDGs may exhibit a deficiency in dark matter content (e.g., van Dokkum et al., 2018; Mancera Piña et al., 2019).
+As for the stellar-to-halo mass relation ($M_{\rm*}-M_{\rm h}$), while Prole et al.
+(2019) propose that LSBGs form a continuous extension of typical dwarfs in the $M_{\rm*}-M_{\rm h}$ diagram, other studies reveal significant deviations for UDGs from the typical $M_{\rm*}-M_{\rm h}$ relation (e.g., van Dokkum et al., 2018).
+Additionally, some giant LSBGs appear to be dominated by baryons at their central regions, akin to the HSBG counterparts (e.g., Lelli et al., 2010; Saburova et al., 2021, 2019); the stellar-to-halo mass relation of the giant LSBGs and HSBGs also exhibits a similarity.
 All of these confusing signs of diversity in matter distribution among LSBGs could offer valuable insights into their varied formation and evolution pathways.
@@
 Over the past two decades, many studies have shown that late-type galaxies and dwarf galaxies consistently adhere to a power-law form of BTFR, despite their wide ranges of baryonic mass, luminosity, and size.
-BTFR could be written as,
-$M_{\rm b}\propto v_{\rm rot}^{\beta},$ (1)
-where the logarithmic slope, $\beta$, typically falls within a range from 3 to 4 (e.g.
-McGaugh et al., 2000; Begum et al., 2008; Papastergis et al., 2016; Lelli et al., 2016b; Sales et al., 2017; Goddy et al., 2021).
+BTFR could be written as, $M_{\rm b}\propto v_{\rm rot}^{\beta}$, where the logarithmic slope, $\beta$, typically falls within a range from 3 to 4 (e.g., McGaugh et al., 2000; Verheijen, 2001; Begum et al., 2008; Papastergis et al., 2016; Lelli et al., 2016a; Sales et al., 2017; Ponomareva et al., 2018; Goddy et al., 2023).
 Notably, some studies have also discovered that early-type galaxies may also follow the BTFR of late-type galaxies (den Heijer et al., 2015; Serra et al., 2012).
@@
 Thus, it can serve as a valuable tool to probe the distribution of baryonic matter and dark matter within galaxies.
-Furthermore, simulations have highlighted the significance of feedback processes in shaping the observed BTFR (e.g.
-Governato et al., 2010; Piontek & Steinmetz, 2011; Dutton, 2012), suggesting that investigating the BTFR can serve as a valuable means to constrain the feedback mechanisms at work within galaxies.
-Additionally, BTFR is highly related to other dynamical relationships such as the radial acceleration relation (e.g.
-McGaugh et al., 2016; Lelli et al., 2017).
+Furthermore, simulations have highlighted the significance of feedback processes in shaping the observed BTFR (e.g., Governato et al., 2010; Piontek & Steinmetz, 2011; Dutton, 2012), suggesting that investigating the BTFR can serve as a valuable means to constrain the feedback mechanisms at work within galaxies.
+Additionally, BTFR is highly related to other dynamical relationships such as the radial acceleration relation (e.g., McGaugh et al., 2016; Lelli et al., 2017).
 Exploring the BTFR could also provide insights into new gravitational theories, as a BTFR with $\beta=4$ is well predicted by the MOdified Newtonian Dynamics (MOND, Milgrom, 1983).
@@
 As anticipated, recent studies examining the BTFR of LSBGs also show some potential signs of tension.
-The previous studies of, for instance, Lelli et al.
-(2016b) (see also Zwaan et al., 1995), have explored that the residuals of BTFR do not change with galactic surface brightness, while some of the recent investigations (Hu et al., 2023; Karunakaran et al., 2020; Rong et al., 2024) reveal that UDGs may significantly deviate from the BTFR of typical late-type galaxies.
+Previous studies have found that the BTFR does not depend on galactic surface brightness or other galactic properties(e.g., Lelli et al., 2016a; Ponomareva et al., 2018; see also Zwaan et al., 1995).
+However, some recent investigations (e.g., Mancera Piña et al., 2019; Guo et al., 2020; Hu et al., 2023; Karunakaran et al., 2020; Rong et al., 2024) reveal that, the particular subsample of LSBGs with large disk sizes, i.e., UDGs, may significantly deviate from the BTFR of typical late-type galaxies, indicative of the possible dependence on size (Mancera Piña et al., 2020).
 Therefore, a thorough and meticulous examination of the BTFR for LSBGs becomes essential.
@@
 In this work, we will first introduce our sample selection in Section 2, and then describe our methods for calculating $M_{\rm b}$ and $v_{\rm rot}$ in Section 3.
-We will then present our results in Section 4, and finally discuss the potential selection biases, the radial acceleration relation, and the possible formation mechanism of LSBGs in Section 5.
+We will then present our results in Section 4, and finally discuss the potential selection biases and the possible formation mechanism of LSBGs in Section 5.
 The summary is in Section 6.
@@
 […]
-From the parent sample, we first remove LSBGs with $g$-band apparent axis ratios $b/a>0.72$, approximately corresponding to the threshold of inclination angles of $i\lesssim 45^{\circ}$, because that the face-on galaxies may introduce large uncertainties in the measurements of the rotation velocities (see equation (10)).
-We then exclude LSBGs with low $\rm HI$ spectral signal-to-noise ratios ${\rm SNR}<10$, and visually remove LSBGs with suspicious HI spectra.
-Due to the bad spatial resolution of ALFALFA (Durbala et al., 2020, being $3.8^{\prime}\times 3.3^{\prime}$;), if a galaxy has a close companion galaxy, its $\rm HI$ spectrum may be contaminated by the companion.
-To eliminate this possibility, we remove the sample galaxies with neighbouring galaxies within the projected radii of $3.8^{\prime}$ (i.e., approximately the beam size of Arecibo) and radial velocity differences of $\Delta v<500{\rm km/s}$.
-The redshift of a LSBG is estimated from the central velocity of its $\rm HI$ spectrum, while the redshift of the neighbouring galaxies comes from the SpecObj and Photoz databases of SDSS.
+$\mu_{0}(B)=\mu_{0}(g)+0.47\times[\mu_{0}(g)-\mu_{0}(r)]+0.17,$ (1)
+to select LSBGs.
+The authors finally selected a sample of LSBGs with the $B$-band apparent central surface brightness of $22.5<\mu_{0}(B)<27.64\ {\rm mag/arcsec^{2}}$, which is treated as the parent sample and used in this work to study the BTFR of LSBGs.
+From the parent sample, we first remove LSBGs with $g$-band apparent axis ratios $b/a>0.72$, approximately corresponding to the threshold of inclination angles of $i\lesssim 45^{\circ}$, because the face-on galaxies may introduce large uncertainties in the measurements of the rotation velocities (see equation (7)).
+We then exclude LSBGs with low $\rm HI$ spectral signal-to-noise ratios ${\rm SNR}<10$ (see Haynes et al., 2018 for the definition of SNR).
+Furthermore, we exclude LSBGs with suspicious HI profiles that indicate potential interactions with other galaxies, serious asymmetric profiles, or profiles with data points affected by poor data quality, as illustrated in Fig.
+1.
+The first two scenarios may suggest an in-equilibrium state of the galaxy, leading to inaccurate estimation of the rotation velocity.
+Figure 1: Examples of suspicious HI spectra from Haynes et al.
+(2018).
+$V_{\rm h}$ denotes the heliocentric velocity, and $W_{\rm 50}$ denotes the 50% peak width of an HI spectrum, obtained by Haynes et al.
+(2018).
+Panel (a): a profile showing potential interactions with other galaxies.
+Panel (b): a seriously asymmetric profile.
+Panel (c): a profile containing bad points due to data quality.
+Due to the bad spatial resolution of ALFALFA, if a galaxy has a close companion galaxy, its $\rm HI$ spectrum may be contaminated by the companion.
+To eliminate this possibility, we remove the galaxies with companions within the projected radii of $3.8^{\prime}$ (i.e., approximately the beam size of Arecibo) and radial velocity differences of $\Delta v<500{\rm km/s}$.
+The redshift of a LSBG is estimated from the central velocity of its $\rm HI$ spectrum  [^1] While the distances of our sample galaxies are obtained directly from Haynes et al., 2018, please refer to their work for more details.
+, while the redshift of the neighbouring galaxies comes from the SpecObj and Photoz databases of SDSS.
 Finally, 204 LSBGs are left in our sample.
-We also calculate the kurtosis coefficients, $k_{\rm 4}$, of the HI spectra (Papastergis et al., 2016; El-Badry et al., 2018) for our sample galaxies, as
+We also calculate the kurtosis coefficient, $k_{\rm 4}$, of the HI spectrum (Papastergis et al., 2016; El-Badry et al., 2018) for each galaxy, as
 $k_{4}=M_{4}/(\sigma^{2})^{2}-3,$ (2)
@@
 […]
-In equations (3) and (4), $\nu$ and $F(\nu)$ represent the frequency and the flux density, respectively, and $\bar{\nu}$ is defined as,
+Here, $\nu$ and $F(\nu)$ represent the frequency and corresponding flux density, respectively. $\bar{\nu}$ is defined as,
 […]
-Here, $\nu_{\rm min}$ and $\nu_{\rm max}$ correspond to the minimal and maximum frequencies between which $W_{20}$ (see Section 3 for the definition of $W_{20}$) is measured, respectively.
-The coefficient $k_{\rm 4}$ quantifies the “peakiness” of a spectrum and thus can indicate whether a spectrum is single-peaked or double-horned.
-The single-peaked galaxies may have kinematics strongly influenced by velocity dispersion (El-Badry et al., 2018), or HI disks that are not extended enough to trace the flat part of the rotation curve (Papastergis et al., 2016).
-As a consequence, their HI-widths might not represent their rotation velocities (McGaugh, 2012; Verheijen, 1997).
-Therefore, we perform a cut at the HI spectrum, $k_{\rm 4}<-1$, in order to remove the single-peaked galaxies.
+where $\nu_{\rm min}$ and $\nu_{\rm max}$ correspond to the minimal and maximum frequencies between which $W_{20}$ (see Section 3 for the definition of $W_{20}$) is measured, respectively.
+$k_{\rm 4}$ can distinguish the single-horned and double-horned spectra (see, e.g., Fig.
+1 of El-Badry et al., 2018).
+The galaxies with single-horned HI lines may have kinematics strongly influenced by velocity dispersion (El-Badry et al., 2018) or beam smearing, or HI disks that are not extended enough to trace the flat part of the rotation curve (Papastergis et al., 2016).
+As a consequence, their HI-widths cannot represent the rotation velocities (McGaugh, 2012; Verheijen, 1997).
+Therefore, we perform a cut at the HI spectrum, $k_{\rm 4}<-1$, in order to remove the galaxies with single-horned HI lines (see Fig.
+3 of El-Badry et al., 2018 for the threshold).
 […]
+Figure 2: A comparison of properties of selected LSBGs and HSBGs.
+The green-filled circles represent LSBGs (the dark-green and light-green depict the galaxies with apparent axis ratios $b/a>0.3$ and $b/a\leq 0.3$, respectively), while the grey-filled circles represent HSBGs.
+Panel (a): axis ratio $b/a$ versus mean surface density $\braket{\Sigma}_{*,\rm{eff}}$.
+Panel (b): $g-i$ color versus gas fraction $f_{\rm{g}}$.
+Panel (c): scale length $R_{\rm{d}}$ versus baryonic mass $M_{\rm{b}}$.
+Panel (d): $v_{\rm rot}$ versus $f_{\rm b}(R_{\rm HI})$.
+Please refer to Section 2 and Section 5.2 for the definitions of these parameters.
+The black-solid line and the blue-shaded region in panel d represent the best-fitting linear line and the bin used for performing the two-sample Kuiper tests, respectively, as described in Section 5.2.
 3 Method
-We study the BTFRs of LSBGs and HSBGs by directly estimating the baryonic masses, $M_{\rm b}$, and rotation velocities, $v_{\rm rot}$, of the galaxies in the two samples.
-The baryonic mass of a galaxy in our samples is mainly composed of the mass of stars and the mass of cold gas.
+The baryonic mass of a galaxy is mainly composed of the stellar mass and gas mass.
 The stellar mass $M_{\rm*}$ can be estimated with the galaxy luminosity in the $W1$-band (denoted as $L_{W1}$) from the Wide-field Infrared Survey Explorer (Wright et al., 2010), using the method proposed by McGaugh & Schombert (2015), as
 $M_{\rm*}/M_{\rm\odot}=0.45L_{W1}/L_{\rm\odot}.$ (6)
-As for the mass of the cold gas, $M_{\rm g}$, we estimate it by adding up the masses of HI and helium,
-$M_{\rm g}=1.33M_{\rm HI},$ (7)
-where $M_{\rm HI}$ is the HI mass.
-Equation (7) will certainly underestimate the mass of the gas since it neglects the mass of the molecular gas and hot ionized gas.
-However, the molecular gas fraction in the total gas component is low and thus negligible (McGaugh, 2012).
-Also, the hot ionized gas usually resides outside the HI disk where our rotational velocity is measured (Dutton, 2012), so it can be neglected.
-Therefore, approximately we estimate the baryonic mass of a galaxy as,
-$M_{\rm b}=M_{\rm*}+M_{\rm g}.$ (8)
-For the rotation velocity, we estimate it using the $20\%$ peak width of a HI spectrum (Guo et al., 2020; Hu et al., 2023; Rong et al., 2024, $W_{\rm 20}$;), as,
+As for the gas mass, $M_{\rm g}$, we estimate it by adding up the masses of HI and helium, where $M_{\rm HI}$ is the HI mass (see Haynes et al., 2018 for the method of estimating $M_{\rm HI}$).
+The baryonic mass of a galaxy is then estimated as $M_{\rm b}=M_{\rm*}+M_{\rm g}$.
+Note that $M_{\rm b}$ does not include the hot ionized gas and molecular gas, as their fractions are expected to be small compared to the HI fraction (McGaugh, 2012; Zhong et al., 2023), and thus are negligible.
+We use the $20\%$ peak width of an HI spectrum (Guo et al., 2020; Hu et al., 2023; Rong et al., 2024, $W_{\rm 20}$;; see the Method section of Guo et al.
+(2020) for the method of estimation of $W_{\rm 20}$)to calculate the rotation velocity $v_{\rm rot}$.
+Guo et al.
+(2020) have demonstrated that $W_{\rm 20}$ can accurately trace galactic $v_{\rm rot}$, which can be estimated as,
 $v_{\rm rot}=\frac{W_{20}}{2\sin i},$ (7)
 where $i$ is the inclination angle.
-$W_{20}$ employed in equation (9) has been corrected for instrumental broadening and redshift (Guo et al., 2020; Kent et al., 2008).
-In equation (9), we assume an alignment between the inclination angle of the stellar disk and HI disk in a galaxy, and then use the inclination angle of the stellar disk to estimate the rotation velocity.
-The inclination is estimated from the apparent axis ratio, as (e.g.
-Guo et al., 2020; Hu et al., 2023; Du et al., 2019; Rong et al., 2024; Begum et al., 2008),
-$\sin i$ $=\sqrt{\frac{1-(b/a)^{2}}{1-q_{0}^{2}}},$ (10)
-where $q_{0}$ denotes the intrinsic axis ratio seen edge-on.
-We take $q_{0}=0.2$, which is commonly used in the previous studies (e.g.
-Du et al., 2019; Guo et al., 2020; Giovanelli et al., 1997; Tully et al., 2009).
-For those edge-on galaxies with $g$-band $b/a\leq 0.2$, we set $\sin i=1$.
-Considering that the stellar disk and gas disk in galaxies may not be perfectly co-planar and often exhibit a small inclination difference of ${\rm\delta}i<20^{\circ}$ (e.g.
-Gault et al., 2021), we assume that the misalignment angle between the gas disk and the stellar disk, ${\rm\delta}i$, follows a Gaussian distribution centred at $0^{\circ}$ with a standard deviation of $\sigma=20^{\circ}$.
-We treat $\sigma=20^{\circ}$ as the uncertainty associated with $i$ in our study.
-4 Results
-In Figure 1, we observe that the distributions of the LSBGs and HSBGs exhibit no significant deviation in the $v_{\rm rot}-M_{\rm b}$ diagram. Subsequently, we impose a power-law equation,
+$W_{20}$ employed in equation (7) has been corrected for instrumental broadening and redshift (Guo et al., 2020; Catinella et al., 2012; Kent et al., 2008).
+The errors of $W_{\rm 20}$ are estimated following the method described by Guo et al.
+(2020).
+Please refer to the Method section of their paper for details.
+In equation (7), we assume an alignment between the inclination angles of the stellar disk and HI disk in a galaxy and then use the inclination angle of the stellar disk to estimate the rotation velocity.
+The inclination is estimated from the $g-$band apparent axis ratio, as (e.g., Guo et al., 2020; Hu et al., 2023; Du et al., 2019; Rong et al., 2024; Begum et al., 2008), $\sin i=\sqrt{\frac{1-(b/a)^{2}}{1-q_{0}^{2}}}$, where $q_{0}$ denotes the intrinsic axis ratio seen edge-on.
+We take $q_{0}=0.2$, which is commonly used in the previous studies (e.g., Du et al., 2019; Guo et al., 2020; Giovanelli et al., 1997; Tully et al., 2009).
+For those edge-on galaxies with $b/a\leq q_{0}$, we set $\sin i=1$.
+Previous studies have revealed that the stellar disk and gas disk in galaxies may not be perfectly co-planar, but often exhibit a small inclination difference of ${\rm\delta}i<20^{\circ}$ (e.g., Starkenburg et al., 2019; Guo et al., 2020; Gault et al., 2021).
+In order to take this misalignment into account, we assume ${\rm\delta}i$ following a Gaussian distribution centred at $0^{\circ}$ with a standard deviation of $\sigma_{i}=20^{\circ}$.
+We treat $\sigma_{i}=20^{\circ}$ as the uncertainty associated with $i$ in our study.
+Note that we do not apply corrections for the velocity dispersion in equation (7), since the effects of the dispersion are severe under $v_{\rm rot}\lesssim 40\ {\rm km/s}$ (summarized by Lelli, 2022) while $v_{\rm rot}$ of our sample galaxies are well above $\sim 50\ {\rm km/s}$.
+4 BTFR
+In Fig. 3, we observe that the distributions of LSBGs and HSBGs exhibit no significant deviation in the $v_{\rm rot}-M_{\rm b}$ diagram. We impose a power-law equation,
 $\log(M_{\rm b})=\alpha+\beta\log(v_{\rm rot}),$ (8)
-where $\alpha$ and $\beta$ are both free parameters, to fit the BTFRs of the HSBGs and LSBGs, respectively.
-The best-fitting parameters of the BTFR of HSBGs are $\alpha_{\rm HSBG}=1.83\pm 0.51$ and $\beta_{\rm HSBG}=3.91\pm 0.24$, while the best-fitting parameters of LSBGs are $\alpha_{\rm LSBG}=-1.45\pm 1.49$ and $\beta_{\rm LSBG}=5.67\pm 0.78$.
+where $\alpha$ and $\beta$ are free parameters to fit the BTFRs of the HSBGs and LSBGs using scipy.odr package, respectively.
+To estimate the errors of $\alpha$ and $\beta$, we apply the Monte Carlo method.
+We generate data samples based on the original ($v_{\rm rot}$, $M_{\rm b}$) values and their associated uncertainties.
+We assume Gaussian distributions for $v_{\rm rot}$ and $M_{\rm b}$ centered at their origins, with the $1\sigma$ regions matching the error bars.
+This process yields a set of ($\alpha$, $\beta$) values.
+We repeat this sampling procedure 5,000 times to obtain 5,000 sets of ($\alpha$, $\beta$) values.
+The standard deviations of these parameters ($\sigma_{x,{\rm MC}}$) are then combined with the fitting errors ($\sigma_{x,{\rm odr}}$) to determine the uncertainties of $\alpha$ and $\beta$, as
+$\sigma_{x}=\sqrt{\sigma_{x,{\rm MC}}^{2}+\sigma_{x,{\rm odr}}^{2}},$ (9)
+where the subscript $x$ denotes $\alpha$ or $\beta$.
+The best-fitting parameters of the BTFR are $\alpha_{\rm HSBG}=1.83\pm 0.51$ and $\beta_{\rm HSBG}=3.91\pm 0.24$ for HSBGs, while are $\alpha_{\rm LSBG}=-1.45\pm 1.49$ and $\beta_{\rm LSBG}=5.67\pm 0.78$ for LSBGs.
 We further split LSBGs and HSBGs into the different $\log v_{\rm{rot}}$ bins.
 Within each bin, we calculate the median value of $\log M_{\rm{b}}$ and the corresponding $1\sigma$ standard deviation for both samples.
-As shown in Figure 1, the binned points of LSBGs (depicted by the blue squares), as well as the binned points of HSBGs (depicted as the black squares), comfortably lie within the $1\sigma$ uncertainty region of the tightest BTFR of typical late-type galaxies reported by Lelli et al.
+As shown in Fig.
+3, the binned points of LSBGs (depicted by the blue squares) and HSBGs (depicted as the black squares) comfortably lie within the $1\sigma$ uncertainty region of the tightest BTFR of typical late-type galaxies reported by Lelli et al.
 (2019) (the red-solid line and the red-shaded region denote the best-fitting BTFR of typical late-type galaxies and its $1\sigma$ uncertainty, respectively, as shown in Fig.
-1), with $\beta\simeq 3.85\pm 0.09$.
+3), with $\beta\simeq 3.85\pm 0.09$.
 Therefore, we conclude that the BTFRs of LSBGs and HSBGs show no significant deviation. Both galaxy samples follow the BTFR of typical late-type galaxies, with a BTFR slope of approximately 4.
-Figure 1: $v_{\rm rot}-M_{\rm b}$ diagram for our samples.
-The green circles represent LSBGs, while the grey circles represent HSBGs, and the orange circles are SPARC galaxies from Lelli et al.
+Our results are consistent with many previous observational (e.g., McGaugh, 2012; Lelli et al., 2016a; Papastergis et al., 2016) and simulation (e.g., Dutton et al., 2017) results.
+However, it is also worth noting that some other studies alternatively report smaller slopes ($\beta\sim 3$, e.g., Goddy et al., 2023; Ponomareva et al., 2018).
+Figure 3: The $v_{\rm rot}-M_{\rm b}$ diagram for our samples.
+The green- and the grey-filled circles are sthe same as those in Figure 2.
+The orange-filled circles are SPARC galaxies from Lelli et al.
 (2019).
 All the error bars denote $\rm 1\sigma$ uncertainties.
-The blue squares represent the median values of LSBGs while the black squares represent that of HSBGs.
+The blue squares represent the median values of LSBGs while the black squares represent those of HSBGs.
 The red-solid line is the tightest BTFR reported by Lelli et al.
@@
 The red-shaded area is the $\rm 1\sigma$ intrinsic scatter of the red-solid line, which is considered as the $\rm 1\sigma$ uncertainty of the tightest BTFR.
-The sub-panels display the number-density distribution of LSBGs (green histograms) and HSBGs (grey histograms).
 5 Discussion
-5.1 Selection biases
-In our study, the SNR threshold used to select the LSBG and HSBG samples may exclude some galaxies with similar HI masses but wider HI line widths from our parent sample, which might potentially introduce deviations in the slope of BTFR.
-To investigate the possible selection effects, we adjust the SNR threshold by increasing it from 10 to 15, or 20.
+5.1 Potential biases
+In our study, the SNR threshold used to select the LSBG and HSBG samples may exclude some galaxies with similar HI masses but wider HI line widths, potentially introducing deviations in the slope of the BTFR.
+To investigate the possible selection effects, we adjust the SNR threshold by increasing it from 10 to 15, and then to 20.
 We find that, for HSBGs, the best-fitting $\beta$ of their BTFR changes to $4.05\pm 0.21$ for SNR $>15$, and $4.05\pm 0.26$ for SNR $>20$, still in good agreement with the best-fitting $\beta$ value for SNR $>10$.
-For LSBGs, the best-fitting $\beta$ becomes $\beta\sim 6.04\pm 0.66$ (SNR $>15$) and $\sim 6.13\pm 0.82$ (SNR $>20$), which is also similar to the $\beta\sim 5.67\pm 0.78$ when applying the selection criterion of SNR$>10$, considering the large uncertainties of the $\beta$ values.
-Additionally, the binned points of LSBGs and HSBGs within the different $\log v_{\rm{rot}}$ for the different SNR thresholds are always located within the $1\sigma$ uncertainty region of the BTFR of typical late-type galaxies.
-Consequently, we conclude that, for both of LSBG and HSBG samples, the selection effect introduced by the SNR threshold is negligible.
-5.2 The radial acceleration relation
-Apart from BTFR, recently, a series of studies (McGaugh et al., 2016; Lelli et al., 2017; Li et al., 2018; Richards et al., 2018; Rong et al., 2018; Chae et al., 2020) have revealed another tight correlation known as the radial acceleration relation (RAR).
-RAR indicates that the acceleration $g_{\rm bar}$ produced by the baryonic mass and the acceleration $g_{\rm obs}$ produced by the dynamical mass at any radius in many types of galaxies consistently follow a correlation of
-$g_{\rm obs}=\frac{g_{\rm bar}}{1-e^{-\sqrt{g_{\rm bar}/a_{0}}}},$ (12)
-where $a_{0}\simeq(1.20\pm 0.02\pm 0.24)\times 10^{-10}\ {\rm m/s^{2}}$ (McGaugh et al., 2016; Lelli et al., 2017) is constant.
-In the $\rm\Lambda CDM$ paradigm, RAR plausibly suggests a universally established link between the mass distributions of baryonic matter and dark matter in galaxies.
-We estimate $g_{\rm obs}$ and $g_{\rm bar}$ of each galaxy at its HI-radius ($R_{\rm HI}$) using the following equations,
-$g_{\rm obs}=\frac{v_{\rm rot}^{2}}{R_{\rm HI}},$ (13)
-$g_{\rm bar}=x_{\rm*}\frac{GM_{\rm*}}{R^{2}_{\rm HI}}+x_{\rm g}\frac{GM_{\rm g}}{R^{2}_{\rm HI}},$ (14)
-where $G$ is the gravitational constant.
-$x_{\rm*}$ and $x_{\rm g}$ in equation (14) are correction factors assuming that the cold gas and stars are distributed in two razor-thin disks, respectively.
-They are calculated as (Freeman, 1970),
-$x_{\rm*}={\rm I_{0}}(\frac{R_{\rm HI}}{2R_{d,*}}){\rm K_{0}}(\frac{R_{\rm HI}}{2R_{d,*}})-{\rm I_{1}}(\frac{R_{\rm HI}}{2R_{d,*}}){\rm K_{1}}(\frac{R_{\rm HI}}{2R_{d,*}}),$ (15)
-$x_{\rm g}={\rm I_{0}}(\frac{R_{\rm HI}}{2R_{d,g}}){\rm K_{0}}(\frac{R_{\rm HI}}{2R_{d,g}})-{\rm I_{1}}(\frac{R_{\rm HI}}{2R_{d,g}}){\rm K_{1}}(\frac{R_{\rm HI}}{2R_{d,g}}),$ (16)
-where $\rm I_{n}$ and $\rm K_{n}$ are the modified Bessel functions of the first and second kind, respectively; $R_{\rm d,*}$ and $R_{\rm g,*}$ involved are the scale lengths of the stellar and HI disks which can be derived by using the same method described in the work of Rong et al.
-(2024).
-$R_{\rm HI}$ corresponds to the outer region of a galaxy where the surface density of HI equals $1\ M_{\rm\odot}/{\rm pc^{2}}$ (Broeils & Rhee, 1997).
-$R_{\rm HI}$ can be estimated from the following tight correlation between HI mass and $R_{\rm HI}$, which has been observed in various types of galaxies including LSBGs (Wang et al., 2016; Gault et al., 2021; Lutz et al., 2018),
-$\log(\frac{2R_{\rm HI}}{\rm kpc})=(0.506\pm 0.003)\log(\frac{M_{\rm HI}}{M_{\rm\odot}})-(3.293\pm 0.009).$ (17)
-In Figure 2, at $R_{\rm HI}$, both our sample LSBGs and HSBGs follow the universal RAR (McGaugh et al., 2016; Lelli et al., 2017, as illustrated by the red components in Fig.
-2).
-The similarity of RARs between LSBG and HSBGs may suggest comparable mass distributions in the two galaxy samples.
-It is worth noting that, at a large radius of a galaxy where $g_{\rm bar}\ll a_{0}$, such as $R_{\rm{HI}}$, the RAR described by equation (12) is actually equivalent to the BTFR with $\beta=4$ (Lelli et al., 2017).
-Figure 2: The RAR of our sample galaxies, overplotted by data from Lelli et al.
-(2016a).
-The circles and squares are the same as those in Figure 1.
-The red-solid line is the best-fitted RAR in previous studies (McGaugh et al., 2016; Lelli et al., 2017), and the grey-dot-dashed line is the identical line.
+For LSBGs, the best-fitting $\beta$ becomes $\sim 6.04\pm 0.66$ (SNR $>15$) and $\sim 6.13\pm 0.82$ (SNR $>20$), which is also similar to the $\beta\sim 5.67\pm 0.78$ when applying the selection criterion of SNR $>10$, considering the large uncertainties of the $\beta$ values.
+Additionally, the binned points of LSBGs and HSBGs within different $\log v_{\rm{rot}}$ bins for the different SNR thresholds are always located within the $1\sigma$ uncertainty region of the BTFR of typical late-type galaxies.
+Consequently, we conclude that, for both the LSBG and HSBG samples, the selection effect introduced by the SNR threshold is negligible.
+It should also be noted that, due to the sensitivity limitations of the Arecibo telescope, galaxies with higher rotational velocities $v_{\rm rot}$ exhibit lower signal-to-noise ratios in their HI spectra under a fixed total HI flux, leading to a bias in our parent sample (Haynes et al., 2011; Dou et al., 2024).
+Another potential bias may arise from the $k_{4}$ cut.
+Some rotation-supported galaxies with small $v_{\rm rot}$ may exhibit single-horned spectra due to beam-smearing effects (Swaters et al., 2003; Di Teodoro & Fraternali, 2015), leading to their exclusion by the $k_{4}$ cut.
+Additionally, galaxies with larger inclination angles tend to have smaller $k_{4}$ values (El-Badry et al., 2018).
+We study the effect of the $k_{4}$ cut by adjusting the $k_{4}$ threshold from -1.0 to -1.2, and then from -1.0 to -0.8.
+The BTFR slope of LSBGs changes from $\sim 5.67\pm 0.78$ ($k_{4}<-1.0$) to $\sim 5.38\pm 0.70$ ($k_{4}<-1.2$) and $\sim 4.28\pm 0.32$ ($k_{4}<-0.8$), while the slope of HSBGs changes from $\sim 3.91\pm 0.24$ ($k_{4}<-1.0$) to $\sim 3.78\pm 0.24$ ($k_{4}<-1.2$) and $\sim 3.85\pm 0.16$ ($k_{4}<-0.8$), respectively.
+It is worth noting that when the $k_{4}$ threshold exceeds $-0.8$, some galaxies appear to deviate from the BTFR.
+However, these galaxies may be strongly dominated by velocity dispersions, as they typically have unreliable $v_{\rm rot}\lesssim 40\ {\rm km/s}$ (Lelli, 2022; El-Badry et al., 2018).
+Since we cannot determine whether the single-horned HI spectra of these galaxies are due to the beam-smearing effect or the dominance of the velocity dispersions, we keep $k_{4}<-1$ to exclude these galaxies.
+Additionally, possible bias emerges from using $W_{20}$ or $W_{50}$ (the $50\%$ peak width of an HI spectrum) to estimate the rotation velocities.
+However, we find that the average difference between the $\log v_{\rm rot}$ estimated from $W_{20}$ and $W_{50}$ is quite small, even smaller than the uncertainties of $\log v_{\rm rot}$.
+Therefore, using $W_{20}$ or $W_{50}$ would not significantly affect our results.
+Besides, we also test whether using the different methods of estimating $M_{\rm*}$ (Taylor et al., 2011) or considering the intrinsic thickness $q_{0}$ dependence on $M_{\rm*}$ (Lelli et al., 2016b) would lead to the changes of BTFR slopes of LSBGs.
+We observe no strong changes in BTFR slopes in these tests.
+Furthermore, we notice that the surface brightness limit determined by the optical survey may lead to large uncertainties in $b/a$ estimation for LSBGs (Mancera Piña et al., 2024).
+However, the large uncertainties should result in a larger scatter, rather than a systematic bias or a change of BTFR slope.
 5.2 Implications on the formation mechanisms of LSBGs
-Previous theoretical studies posit that LSBGs may form in low-density dark matter halos (e.g.
-Dekel & Silk, 1986; McGaugh, 1992).
+Previous theoretical studies posit that LSBGs may form in low-density dark matter halos (e.g., Dekel & Silk, 1986; McGaugh, 1992).
 However, this model predicts a significant difference in the BTFRs between LSBGs and HSBGs (McGaugh, 2021), which contradicts our findings.
 Therefore, the formation model based on low-density dark matter halos may not refer to LSBGs.
-Many observational studies also suggest that LSBGs might be dominated by dark matter compared with HSBGs (e.g.
-de Blok & McGaugh, 1997).
+Many observational studies also suggest that LSBGs might be dominated by dark matter compared with HSBGs (e.g., de Blok & McGaugh, 1997; Swaters et al., 2003; Mancera Piña et al., 2022; Read et al., 2017).
 Therefore, we delve deeper into understanding the fraction of dark matter in LSBGs.
-Note that BTFR could be expressed as  [^1] Equation (19) is from the lecture on Tully-Fisher relation; see also Aaronson et al.
+Note that BTFR could be expressed as  [^2] Equation (10) is from the lecture on the Tully-Fisher relation; see also Aaronson et al.
 (1979) and Zwaan et al.
@@
 $M_{\rm b}\simeq\frac{f_{\rm b}^{2}(R)}{Gg_{\rm bar}(R)}v_{\rm rot}^{4}.$ (10)
-where $R$ corresponds to a large radius in the outer region of a galaxy (such as $R_{\rm{HI}}$), and
-$f_{\rm b}(R)=\frac{M_{\rm b}}{M_{\rm b}+M_{\rm DM}(R)},$ (19)
-is the baryonic mass fraction within $R$, where $M_{\rm{DM}}(R)$ denotes the dark matter mass enclosed within $R$.
-$g_{\rm bar}(R)$ is the baryonic acceleration at $R$, which could be calculated with equation (14).
-Consequently, $f_{\rm b}(R_{\rm{HI}})$ of the LSBGs and HSBGs can be estimated with equation (18).
-We examine the variation of $f_{\rm b}(R_{\rm HI})$ as a function of $v_{\rm{rot}}$ for both LSBGs and HSBGs.
-As shown in Fig.
-3, in statistics, $f_{\rm b}(R_{\rm HI})$ increases with rising $v_{\rm{rot}}$, with a slope of $\sim 1.02$ using a bisector fitting (Isobe et al., 1990).
-Since the dark matter halo mass of a galaxy is tightly related to $v_{\rm{rot}}$, as $M_{\rm{h}}\propto v_{\rm{rot}}^{3}$ (Mo et al., 1998), the correlation between $f_{\rm b}(R_{\rm HI})$ and $v_{\rm{rot}}$ hints at a positive association between the baryonic mass fractions and dark matter halo masses for these galaxies (Ayromlou et al., 2023; Duffy et al., 2010).
+Here, the baryonic mass fraction within a radius $R$ is defined as $f_{\rm b}(R)=M_{\rm b}/M_{\rm tot}(R)$, where $M_{\rm{tot}}(R)$ denotes the total mass enclosed within $R$.
+$g_{\rm bar}(R)$ is the baryonic acceleration at $R$.
+$g_{\rm bar}(R)$ is calculated by assuming the stellar and gas components in a galaxy follow the razor-thin disk distribution, using the method outlined in Read et al.
+(2017) (see equation (5) of Read et al.
+(2017) for details).
+Note that equation (10) holds true only when $R$ is large.
+Here, we utilize the HI radius, $R_{\rm HI}$, defined as the radius where the HI surface density equals $1\ M_{\rm\odot}/{\rm pc^{2}}$.
+$R_{\rm HI}$ can be estimated based on the strong correlation between $R_{\rm HI}$ and $M_{\rm HI}$ (Wang et al., 2016; Gault et al., 2021; Lutz et al., 2018),
+$\log\left(\frac{2R_{\rm HI}}{\rm kpc}\right)=(0.506\pm 0.003)\log\left(\frac{M_{\rm HI}}{M_{\rm\odot}}\right)-(3.293\pm 0.009).$ (11)
+Consequently, we calculate $f_{\rm b}(R_{\rm HI})$ from equation (10) and then analyze the $f_{\rm b}(R_{\rm HI})\--v_{\rm rot}$ relationship for the LSBGs and HSBGs.
+As shown in panel (d) of Figure 2, in statistics, $f_{\rm b}(R_{\rm HI})$ increases with rising $v_{\rm{rot}}$, with a slope of $\sim 1.02$ using a bisector fitting (Isobe et al., 1990).
+Since the dark matter halo mass ($M_{\rm h}$) of a galaxy is tightly related to $v_{\rm{rot}}$, as $M_{\rm{h}}\propto v_{\rm{rot}}^{3}$ (Mo et al., 1998), the correlation between $f_{\rm b}(R_{\rm HI})$ and $v_{\rm{rot}}$ hints at a positive association between the baryonic mass fractions and dark matter halo masses for these galaxies (Ayromlou et al., 2023; Duffy et al., 2010).
+Similar trends have also been observed based on spatially resolved data (e.g., Lelli et al., 2017; Posti et al., 2019; Mancera Piña et al., 2022; Di Teodoro et al., 2023).
 This correlation does not exceed our expectations, given that our galaxy samples are dominated by the low-mass galaxies.
-Halos with lower $M_{\rm{h}}$ have shallower potentials (Duffy et al., 2010), making it more difficult to prevent the galactic interstellar medium (ISM) against expulsion by the feedback processes in these shallower potentials, which leads to lower $f_{\rm b}(R_{\rm{HI}})$ (e.g.
-Bryan et al., 2013; Duffy et al., 2010).
-Subsequently, in order to compare the $f_{\rm b}(R_{\rm{HI}})$ of the LSBG and HSBG samples with similar dark matter halo masses, we further focus on a specific $v_{\rm{rot}}$ bin defined by $\log v_{\rm rot}\in[1.87,2.05]$ (as depicted by the blue-shaded region in Fig.
-3).
+Halos with lower $M_{\rm{h}}$ have shallower potentials (Duffy et al., 2010), making it more difficult to prevent the galactic interstellar medium (ISM) against expulsion by the feedback processes in these shallower potentials, which leads to lower $f_{\rm b}(R_{\rm{HI}})$ (e.g., Bryan et al., 2013; Duffy et al., 2010).
+Subsequently, in order to compare the $f_{\rm b}(R_{\rm{HI}})$ of the LSBG and HSBG samples with similar dark matter halo masses, we further focus on a specific $v_{\rm{rot}}$ bin defined by $\log v_{\rm rot}\in[1.87,2.05]$ (as depicted by the blue-shaded region in panel (d) of Fig.
+2).
 We have 52 LSBGs and 52 HSBGs in this small bin, exhibiting similar distributions of $v_{\rm rot}$ with median $\log v_{\rm rot}$ values of $1.94\pm 0.05$ (for LSBGs) and $1.97\pm 0.05$ (for HSBGs).
@@
 The median values of the $f_{\rm b}(R_{\rm HI})$ distributions of LSBGs and HSBGs are $-0.94\pm 0.15$ and $-0.91\pm 0.30$, respectively; the Kuiper test for the two $f_{\rm b}(R_{\rm HI})$ distributions also yields a large $p$-value of $0.34$.
-Therefore, we note that, for the LSBGs and HSBGs with similar halo masses, their baryonic fractions are also comparable.
-LSBGs are not particularly dark matter dominated galaxies, compared with the HSBG counterparts.
+Therefore, we note that, for LSBGs and HSBGs with similar halo masses, their baryonic fractions are also comparable.
+LSBGs may not be particularly dark matter-dominated galaxies compared to their HSBG counterparts; at least within the regimes of $R_{\rm HI}$, the dark matter fractions of LSBGs and HSBGs are similar.
 Furthermore, this result also plausibly suggests comparable feedback power among the LSBGs and HSBGs with similar halo masses (i.e., similar gravitational potentials).
@@
 In other words, these HI-bearing LSBGs may not originate from stronger or weaker feedback.
-Alternatively, our finding aligns with the simulation results where the LSBGs can be well reproduced by higher dark matter halo spins (Pérez-Montaño et al., 2022).
-Therefore, we conclude that our result may lend support to the theorem that LSBGs may originate from high-spin dark matter halos (Mo et al., 1998).
-Figure 3: The $v_{\rm rot}-f_{\rm b}(R_{\rm HI})$ relation of our sample galaxies.
-The circles are the same as those in Figure 1.
-The blue-shaded area represents the bin in which we perform Kuiper tests in Section 5.3.
-The black-dashed line is the best-fitted line for the data points.
-5.4 Implications on MOND
-Except for the aforementioned formation models of LSBGs in the framework of $\rm\Lambda CDM$, MOND may also offer a reasonable explanation for the BTFR and RAR of LSBGs (e.g.
-Milgrom, 1983; Wittenburg et al., 2020).
-As demonstrated in Section 4 and 5.2, both LSBGs and HSBGs adhere to a BTFR with a slope indistinguishable from 4, and may follow the same RAR as given by equation (12).
-These results adhere to the theoretical predictions by MOND (Milgrom, 1983).
+Alternatively, our finding aligns with the simulation results where the LSBGs can be well reproduced by higher dark matter halo spins (e.g., Pérez-Montaño et al., 2022; Di Cintio et al., 2019).
+Therefore, we conclude that our result may lend support to the theorem that LSBGs may originate from high-spin dark matter halos (e.g., Mo et al., 1998).
+Except for the aforementioned formation models of LSBGs in the framework of $\rm\Lambda CDM$, MOND may also offer a reasonable explanation for the BTFR of LSBGs (e.g., Milgrom, 1983; Wittenburg et al., 2020).
+As demonstrated in Section 4, both LSBGs and HSBGs follow a BTFR with a slope indistinguishable from 4.
+This result adheres to the theoretical predictions by MOND (Milgrom, 1983).
 Therefore, our result cannot exclude the MOND theory.
+Nevertheless, several studies on the dynamics of galaxies have produced results that deviate from MOND’s predictions (e.g., Ren et al., 2019; Mercado et al., 2024; Mancera Piña et al., 2024; Khelashvili et al., 2024).
+Therefore, more compelling studies are needed to better determine if MOND is a promising framework.
 6 Summary
 Taking advantage of the ALFALFA HI survey data, we have selected the HI-bearing LSBGs and HSBGs, and estimated their baryonic masses $M_{\rm b}$ and rotation velocities $v_{\rm rot}$.
-Our findings reveal that, statistically, the HI-bearing LSBGs and HSBGs also conform to the BTFR found in typical late-type galaxies as well as ordinary dwarf galaxies with slope of $\beta\sim 4$.
+Our findings reveal that, statistically, the HI-bearing LSBGs and HSBGs also conform to the BTFR found in typical late-type galaxies as well as ordinary dwarf galaxies with slope $\sim 4$.
 These results hint that LSBGs may owe their properties to the high spins of their host dark matter halos, rather than the halo densities or feedback processes.
 Additionally, our findings do not rule out the Modified Newtonian Dynamics (MOND) theory.
-Our findings conflict with the BTFR results of UDGs (Karunakaran et al., 2020; Hu et al., 2023; Rong et al., 2024).
+Our findings conflict with the BTFR results of UDGs (Mancera Piña et al., 2019; Guo et al., 2020; Hu et al., 2023; Karunakaran et al., 2020; Rong et al., 2024).
 This is primarily because that our LSBG sample contains very few UDGs.
 Indeed, almost all of our LSBGs are classic LSBGs, rather than UDGs with extremely large effective radii.
-Therefore, we propose that UDGs may be a distinct galaxy population with different BTFR and mass distributions, compared with typical LSBGs.
+Therefore, we propose that UDGs may be a distinct galaxy population with different BTFR and mass distributions, compared with classical LSBGs.
 They may also have distinct formation mechanisms that warrant further investigations.
 Acknowledgements
-We sincerely thank Xufen Wu, Haochen Jiang, Jingshuo Yang, Shijiang Chen and Yu Chen for the inspiring discussions during this study.
+We express our sincere thanks to the referee for the detailed comments and suggestions.
+We also thank Xufen Wu, Haochen Jiang, Jingshuo Yang, Shijiang Chen and Yu Chen for the inspiring discussions during this study.
 Y.R.
-acknowledges supports from the NSFC grant 12273037, and the CAS Pioneer Hundred Talents Program (Category B), as well as the USTC Research Funds of the Double First-Class Initiative (No.
-YD2030002013).
+acknowledges supports from the NSFC grant 12273037, the CAS Pioneer Hundred Talents Program (Category B), the USTC Research Funds of the Double First-Class Initiative (No.
+YD2030002013), and the research grants from the China Manned Space Project (the second-stage CSST science projects: “Investigation of small-scale structures in galaxies and forecasting of observations” and “CSST study on specialized galaxies in ultraviolet and multi-band”).
 H.H.

### v4

@@
-The Baryonic Tully-Fisher Relation of HI-bearing Low Surface Brightness Galaxies Implies Their Formation Mechanism
+The baryonic Tully-Fisher relation of HI-bearing low-surface brightness galaxies implies their formation mechanism
 Abstract
-We investigate the baryonic Tully-Fisher relation in low surface brightness galaxies selected from the Arecibo Legacy Fast ALFA survey.
-We find that the $\rm HI$-bearing low surface brightness galaxies still follow the baryonic Tully-Fisher relation of typical late-type galaxies, with a slope of approximately 4 in the baryonic mass versus rotational velocity diagram on the logarithmic scale, i.e., $M_{\rm{b}}\propto v_{\rm{rot}}^{4}$.
-Our findings suggest that the matter distributions in low surface brightness galaxies may resemble that of general late-type galaxies, and hint that low surface brightness galaxies may not originate from dark matter halos of low densities or stronger/weaker feedback processes, but may emerge from dark matter halos with high spin values.
+We investigate the baryonic Tully-Fisher relation in low-surface brightness galaxies selected from the Arecibo Legacy Fast ALFA survey.
+We find that the $\rm HI$-bearing low-surface brightness galaxies still follow the baryonic Tully-Fisher relation of typical late-type galaxies, with a slope of approximately 4 in the baryonic mass versus rotational velocity diagram on the logarithmic scale, i.e., $M_{\rm{b}}\propto v_{\rm{rot}}^{4}$.
+Our findings suggest that the matter distributions in low-surface brightness galaxies may resemble that of general late-type galaxies, and hint that low-surface brightness galaxies may not originate from dark matter halos of low densities or stronger/weaker feedback processes, but may emerge from dark matter halos with high spin values.
 keywords:
 galaxies: kinematics and dynamics – galaxies: formation – galaxies: evolution
-pubyear: 2024pagerange: The Baryonic Tully-Fisher Relation of HI-bearing Low Surface Brightness Galaxies Implies Their Formation MechanismThe Baryonic Tully-Fisher Relation of HI-bearing Low Surface Brightness Galaxies Implies Their Formation Mechanism
+pubyear: 2024pagerange: The baryonic Tully-Fisher relation of HI-bearing low-surface brightness galaxies implies their formation mechanismThe baryonic Tully-Fisher relation of HI-bearing low-surface brightness galaxies implies their formation mechanism
 1 Introduction
-Low surface brightness galaxies (LSBGs) constitute a special population of galaxies with central surface brightness at least one magnitude fainter than that of the sky background (i.e., $B$-band surface brightness $\rm\geq 22.5\ mag\cdot arcsec^{-2}$; Impey & Bothun, 1997; Bothun et al., 1997).
+low-surface brightness galaxies (LSBGs) constitute a special population of galaxies with central surface brightness at least one magnitude fainter than that of the sky background (i.e., $B$-band surface brightness $\rm\geq 22.5\ mag\cdot arcsec^{-2}$; Impey & Bothun, 1997; Bothun et al., 1997).
 Many isolated LSBGs have been observed to exhibit higher HI fractions (e.g., Catinella et al., 2018; Schombert & McGaugh, 2021; Du et al., 2015; He et al., 2020).
-However, some studies (e.g., Martin et al., 2019; Galaz et al., 2002) alternatively found that some LSBGs exhibit the same or even smaller gas fractions compared to high surface brightness galaxies (HSBGs).
+However, some studies (e.g., Martin et al., 2019; Galaz et al., 2002) alternatively found that some LSBGs exhibit the same or even smaller gas fractions compared to high-surface brightness galaxies (HSBGs).
 LSBGs usually have low star formation rates (Wyder et al., 2009; Rong et al., 2020b), low metallicities (Kuzio de Naray et al., 2004; Schombert & McGaugh, 2021), and tend to appear in low-density environment (Pérez-Montaño & Cervantes Sodi, 2019; Mo et al., 1994; Galaz et al., 2011).
@@
 Over the past two decades, many studies have shown that late-type galaxies and dwarf galaxies consistently adhere to a power-law form of BTFR, despite their wide ranges of baryonic mass, luminosity, and size.
-BTFR could be written as, $M_{\rm b}\propto v_{\rm rot}^{\beta}$, where the logarithmic slope, $\beta$, typically falls within a range from 3 to 4 (e.g., McGaugh et al., 2000; Verheijen, 2001; Begum et al., 2008; Papastergis et al., 2016; Lelli et al., 2016a; Sales et al., 2017; Ponomareva et al., 2018; Goddy et al., 2023).
+BTFR could be written as, $M_{\rm b}\propto v_{\rm rot}^{\beta}$, where the logarithmic slope, $\beta$, typically falls within a range from 3 to 4 (e.g., McGaugh et al., 2000; Verheijen, 2001; Begum et al., 2008; Papastergis et al., 2016; Lelli et al., 2016b; Sales et al., 2017; Ponomareva et al., 2018; Goddy et al., 2023).
 Notably, some studies have also discovered that early-type galaxies may also follow the BTFR of late-type galaxies (den Heijer et al., 2015; Serra et al., 2012).
@@
 As anticipated, recent studies examining the BTFR of LSBGs also show some potential signs of tension.
-Previous studies have found that the BTFR does not depend on galactic surface brightness or other galactic properties(e.g., Lelli et al., 2016a; Ponomareva et al., 2018; see also Zwaan et al., 1995).
+Previous studies have found that the BTFR does not depend on galactic surface brightness or other galactic properties(e.g., Lelli et al., 2016b; Ponomareva et al., 2018; see also Zwaan et al., 1995).
 However, some recent investigations (e.g., Mancera Piña et al., 2019; Guo et al., 2020; Hu et al., 2023; Karunakaran et al., 2020; Rong et al., 2024) reveal that, the particular subsample of LSBGs with large disk sizes, i.e., UDGs, may significantly deviate from the BTFR of typical late-type galaxies, indicative of the possible dependence on size (Mancera Piña et al., 2020).
@@
 Therefore, we conclude that the BTFRs of LSBGs and HSBGs show no significant deviation. Both galaxy samples follow the BTFR of typical late-type galaxies, with a BTFR slope of approximately 4.
-Our results are consistent with many previous observational (e.g., McGaugh, 2012; Lelli et al., 2016a; Papastergis et al., 2016) and simulation (e.g., Dutton et al., 2017) results.
+Our results are consistent with many previous observational (e.g., McGaugh, 2012; Lelli et al., 2016b; Papastergis et al., 2016) and simulation (e.g., Dutton et al., 2017) results.
 However, it is also worth noting that some other studies alternatively report smaller slopes ($\beta\sim 3$, e.g., Goddy et al., 2023; Ponomareva et al., 2018).
@@
 The green- and the grey-filled circles are sthe same as those in Figure 2.
-The orange-filled circles are SPARC galaxies from Lelli et al.
+The orange-filled circles are SPARC galaxies (Lelli et al., 2016a) from Lelli et al.
 (2019).
@@
 Therefore, using $W_{20}$ or $W_{50}$ would not significantly affect our results.
-Besides, we also test whether using the different methods of estimating $M_{\rm*}$ (Taylor et al., 2011) or considering the intrinsic thickness $q_{0}$ dependence on $M_{\rm*}$ (Lelli et al., 2016b) would lead to the changes of BTFR slopes of LSBGs.
+Besides, we also test whether using the different methods of estimating $M_{\rm*}$ (Taylor et al., 2011) or considering the intrinsic thickness $q_{0}$ dependence on $M_{\rm*}$ (Lelli et al., 2016c) would lead to the changes of BTFR slopes of LSBGs.
 We observe no strong changes in BTFR slopes in these tests.